Swift Company has a Maintenance Department which provides service to line departments A and B. The cost driver in the Maintenance Department is machine hours worked by the line departments. The Maintenance Department planned variable cost of $8,000 for June. Actual variable cost for the month totaled $15,000. Budget and actual machine hours in the operating departments for the month are shown below. How much variable maintenance cost should be charged to Dept. A for performance evaluation (end of period)?
  Dept. A Dept. B
Budget machine hours 22,000 18,000
Actual machine hours 30,000 20,000
Answer #1

Total budgeted machine hours= 22000+18000= 40000 hours

Planned variable cost per machine hour= Total variable cost/Total budgeted machine hours

= $8000/40000= 0.20 per machine hour

Actual machine hours of Department A= 30000 hours

Variable maintenance cost charged to Department A= Actual machine hours of Department A*Planned variable cost per machine hour

= 30000*0.20= $6000